# The fibering method for singular degenerate Kirchhoff problems with unbalanced-growth operators

### v1

@@ 1. Introduction
{$-m\left(\int_{\Omega}\mathcal{H}(x,|\nabla u|)\,{\rm d}x\right)\mathcal{L}(u)&=\lambda f(u)+g(u)\quad&&\text{in }\Omega,\\ u&>0\quad&&\text{in }\Omega,\\[4.30554pt] u&=0\quad&&\text{on }\partial\Omega,\end{aligned}\right.$ (${\rm P}_{\lambda}$)
@@
First, we mention that hypotheses (H) includes the standard Kirchhoff function $m(s)=a+bs^{\eta}$ with $a,b\in\mathbb{R}^{2}\setminus\{(0,0)\}$, that means we allow degenerate Kirchhoff problems which create the most interesting models in applications. The following operators are included in hypotheses (H), whereby we suppose in all cases that $1<p<N$, $p<q$, and $0\leq\mu(\cdot)\in L^{\infty}(\Omega)$, while we assume $q<p^{*}:=\frac{Np}{N-p}$ for (i) and $q<p^{*}-1$ for (ii)–(iii):
@@
The idea is then to split the Nehari manifold into three disjoint parts minimizing $J$ over two of them to get the required solutions with different energy sign. This method is not new, but it is the first time that it is applied to a very general setting and so no concrete, long calculations are needed. Indeed, we do not only cover the results obtained by Papageorgiou–Repovš–Vetro [27] ($(q,p)$-Laplacian), Papageorgiou–Winkert [26] (weighted $p$-Laplace), Liu–Dai–Papageorgiou–Winkert [20] (double phase operator) or Arora–Fiscella–Mukherjee–Winkert [4] (Kirchhoff double phase operator), but we also have much easier and shorter proofs as in those papers and we also cover new operators within our setting, like the logarithmic double phase operators given in (1.4) and (1.6).
In general, the use of the fibering method along with the Nehari manifold is a very powerful tool and has been further developed by the works of Drábek–Pohozaev [13] and Sun–Wu–Long [28]. Subsequently, several authors have applied this method to various problems of singular type and non-singular type. We refer to works by Alves–Santos–Silva [1] (singular-superlinear Schrödinger equations with indefinite-sign potential), Arora–Fiscella–Mukherjee–Winkert [3] (critical double phase Kirchhoff problems with singular nonlinearity), Chen–Kuo–Wu [7] (Kirchhoff Laplace equations), Fiscella–Mishra [15] (fractional singular Kirchhoff problems), Kumar–Rădulescu–Sreenadh [19] (singular problems with unbalanced growth and critical exponent), Liu–Winkert [21] (double phase problems in $\mathbb{R}^{N}$), Mukherjee–Sreenadh [23] (fractional $p$-Laplace problems), Tang–Cheng [29] (ground state solutions of Nehari-Pohozaev type for Kirchhoff-type problems with general potentials), Wang–Zhao–Zhao [30] (critical Laplace equations with singular term), see also the references therein. For a survey concerning singular problems, we address the reader to the overview article by Guarnotta–Livrea–Marano [17]. It should be mentioned that, in contrast to the results available in the literature (see the list above and also Candito–Guarnotta–Perera [6] and Candito–Guarnotta–Livrea [5]), our method does not require the use of Hardy-Sobolev’s inequality.
The paper is organized as follows. In Section 2 we introduce our function space and recall some basic facts about generalized $N$-functions and related Musielak-Orlicz Sobolev spaces. Further, we prove some auxiliary results and give the precise definition of the Nehari manifold to problem (${\rm P}_{\lambda}$) including its splitting into three disjoint parts. Section 3 discusses some basic estimates which are needed in the sequel while Section 4 gives a detailed study of the Nehari manifold and its properties. Finally, in Section 5, we are able to prove Theorem 1.1.
@@ 2. Preliminaries
Let $M(\Omega)$ be the set of all measurable functions $u\colon\Omega\to\mathbb{R}$. Then, the Musielak-Orlicz space $L^{\varphi}(\Omega)$ is defined by
@@
    Furthermore, if $u\in L^{\varphi}(\Omega)$, then $\rho_{\varphi}(u)<1$ (resp. $=1$; >1)>1)> 1 ) if and only if $\|u\|_{\varphi}<1$ (resp. $=1$; $>1$).
@@
We call a function $\varphi\colon\Omega\times\mathbb{R}\to[0,\infty)$ a generalized $N$-function if $\varphi(\cdot,t)$ is measurable for all $t\in\mathbb{R}$ and $\varphi(x,\cdot)$ is a $N$-function for a.a. $x\in\Omega$. We denote the class of all generalized $N$-functions by $N(\Omega)$.
@@ 1. $\bullet$
Next, we introduce the following functions, useful to compare generalized $N$-functions with suitable power functions. To this end, for given $-\infty<\alpha\leq\beta<+\infty$, we define
@@
$-\infty<i_{k}:=\inf_{s>0}\frac{sk^{\prime}(s)}{k(s)}\leq\sup_{s>0}\frac{sk^{\prime}(s)}{k(s)}=:s_{k}<+\infty.$ (2.5)
@@
It is worth noticing that (2.4) is automatically satisfied when $i_{k}>-1$, due to (2.11). This is the case of the $N$-function $K:=\mathcal{H}(x,\cdot)$ (since (HHL)(H(HL)i) forces $i_{k}>0$) and of the singular term $k:=f$ (see (HHf)).
@@
Proposition 2.6 (applied with $k=\partial_{s}\mathcal{H}(x,\cdot)$), (HHL)(H(HL)i), and (HHL)(H(HL)iii) entail
@@
$\tilde{f}(s):=\begin{cases}f(s)\quad&\text{if }s>0,\\ 0\quad&\text{if }s=0,\\ -f(-s)\quad&\text{if }s<0,\\ \end{cases}\qquad\tilde{g}(s):=\begin{cases}g(s)\quad&\text{if }s>0,\\ 0\quad&\text{if }s=0,\\ -g(-s)\quad&\text{if }s<0.\\ \end{cases}$
@@
$M(s)=\int_{0}^{s}m(t)\,{\rm d}t\quad\text{for all }s\in\mathbb{R},\quad\theta:=\sup_{s>0}\frac{sm(s)}{M(s)},$
@@
$\mathcal{N}^{+}$ :=$:=\{u\in\mathcal{N}\colon\psi_{u}^{\prime\prime}(1)>0\},$
$\mathcal{N}^{0}$ :=$:=\{u\in\mathcal{N}\colon\psi_{u}^{\prime\prime}(1)=0\},$
$\mathcal{N}^{-}$ :=$:=\{u\in\mathcal{N}\colon\psi_{u}^{\prime\prime}(1)<0\}.$
@@
$E^{+}_{u}$ :=$:=\{t\in(0,+\infty)\colon tu\in\mathcal{N}^{+}\},$
$E^{0}_{u}$ :=$:=\{t\in(0,+\infty)\colon tu\in\mathcal{N}^{0}\},$
$E^{-}_{u}$ :=$:=\{t\in(0,+\infty)\colon tu\in\mathcal{N}^{-}\}.$
@@ 3. ${\rm(a_{3})}$
where $\underline{W}_{p}^{q\theta},\overline{W}_{p}^{q\theta}$ are as in (2.3).
@@
Proposition 2.6, (HHL), and the definition of $\theta$ yield
@@ 2. ${\rm(b_{2})}$
In particular, $G$ is a generalized $N$-function satisfying the hypotheses of Proposition 2.8. Moreover, (HHC) and Proposition 2.6 yield
@@ 4. Analysis of the Nehari manifold
with $c,C>0$. Recalling that $1-\gamma_{-}<p$ yields $J(u_{n})\to+\infty$. ∎
@@
Thus, owing to Lemma 3.1, (HHf), Proposition 2.6, and Lemma 3.2, we have
@@
Exploiting Lemma 3.1, (HHg), Proposition 2.6, and Lemma 3.2, we deduce
@@
Fix any $u\in W^{1,\mathcal{H}}_{0}(\Omega)\setminus\{0\}$. We are going to apply Proposition 2.10 to the function $\psi^{\prime}_{u}$. To this end, using Lebesgue’s dominated convergence theorem and Fatou’s lemma, along with (HHm), (HHL), (HHg), (HHf), and the symmetry of $f$, ensure
@@ 5. Proof of the main result
00 0 $\leq\liminf_{n\to\infty}\langle\phi^{\prime}(\overline{t}\nabla u_{n})-\phi^{\prime}(\overline{t}\nabla u),\nabla u_{n}-\nabla u\rangle$
@@
00 0 $\leq\limsup_{n\to\infty}\frac{J(u+t_{n}h)-J(u)}{t_{n}}$ (5.5)
@@
00 0 $\leq\frac{J(\zeta(t_{n}h)(u+t_{n}h))-J(u)}{t_{n}}\leq\frac{J(\zeta(t_{n}h)(u+t_{n}h))-J(\zeta(t_{n}h)u)}{t_{n}}$ (5.6)
@@
00 0 $\leq\limsup_{n\to\infty}\frac{J(\zeta(t_{n}h)(u+t_{n}h))-J(u)}{t_{n}}$
@@
00 0 $\leq\langle T,(u+\varepsilon\varphi)^{+}\rangle$
@@
Notice that |$|\{u+\varepsilon\varphi\leq 0\}|\to 0$ as $\varepsilon\to 0$. Therefore, $\langle T,\varphi\rangle\geq 0\quad\text{for all }\varphi\in W^{1,\mathcal{H}}_{0}(\Omega)$. Since $\varphi$ is arbitrarily chosen, we have $\langle T,\varphi\rangle=0$ for all $\varphi\in W^{1,\mathcal{H}}_{0}(\Omega)$, which means that $u$ is a weak solution to (${\rm P}_{\lambda}$). ∎

### v2

# Degenerate singular Kirchhoff problems in Musielak-Orlicz spaces

@@ 1. Introduction
$\left\{\begin{aligned} -m\left(\int_{\Omega}\mathcal{H}(x,|\nabla u|)\,{\rm d}x\right)\mathcal{L}(u)&=\lambda f(u)+g(u)\quad&&\text{in }\Omega,\\ u&>0\quad&&\text{in }\Omega,\\[4.30554pt] u&=0\quad&&\text{on }\partial\Omega,\end{aligned}\right$. (${\rm P}_{\lambda}$)
@@
First, we mention that hypotheses (H) include the standard Kirchhoff function $m(s)=a+bs^{\eta}$ with $a,b\in\mathbb{R}^{2}\setminus\{(0,0)\}$, that means we allow degenerate Kirchhoff problems which create the most interesting models in applications. The following operators are included in hypotheses (H), whereby we suppose in all cases that $1<p<N$, $p<q$, and $0\leq\mu(\cdot)\in L^{\infty}(\Omega)$, while we assume $q<p^{*}:=\frac{Np}{N-p}$ for (i) and $q+\kappa<p^{*}$ for (ii)–(iii), where $\kappa:=\frac{e}{e+t_{0}}\in(0,1)$ and $t_{0}>0$ is the unique solution of $t=e\log(e+t)$:
@@
The idea is then to split the Nehari manifold into three disjoint parts minimizing $J$ over two of them to get the required solutions with different energy sign. This method is not new, but it is the first time that it is applied to a very general setting and so no concrete, long calculations are needed. Indeed, we do not only cover the results obtained by Papageorgiou–Repovš–Vetro [27] ($(q,p)$-Laplacian), Papageorgiou–Winkert [26] (weighted $p$-Laplacian), Liu–Dai–Papageorgiou–Winkert [20] (double phase operator) or Arora–Fiscella–Mukherjee–Winkert [4] (Kirchhoff double phase operator), but we also have much easier and shorter proofs as in those papers and we also cover new operators within our setting, like the logarithmic double phase operators given in (1.4) and (1.6).
In general, the use of the fibering method along with the Nehari manifold is a very powerful tool and has been further developed by the works of Drábek–Pohozaev [13] and Sun–Wu–Long [28]. Subsequently, several authors have applied this method to various problems of singular type and non-singular type. We refer to works by Alves–Santos–Silva [1] (singular-superlinear Schrödinger equations with indefinite-sign potential), Arora–Fiscella–Mukherjee–Winkert [3] (critical double phase Kirchhoff problems with singular nonlinearity), Chen–Kuo–Wu [7] (Kirchhoff Laplace equations), Fiscella–Mishra [15] (fractional singular Kirchhoff problems), Kumar–Rădulescu–Sreenadh [19] (singular problems with unbalanced growth and critical exponent), Liu–Winkert [21] (double phase problems in $\mathbb{R}^{N}$), Mukherjee–Sreenadh [23] (fractional $p$-Laplace problems), Tang–Cheng [29] (ground state solutions of Nehari–Pohozaev type for Kirchhoff-type problems with general potentials), Wang–Zhao–Zhao [30] (critical Laplace equations with singular term), see also the references therein. For a survey concerning singular problems, we address the reader to the overview article by Guarnotta–Livrea–Marano [17]. It should be mentioned that, in contrast to the results available in the literature (see the list above and also Candito–Guarnotta–Perera [6] and Candito–Guarnotta–Livrea [5]), our method does not require the use of Hardy-Sobolev’s inequality.
The paper is organized as follows. In Section 2 we introduce our function space and recall some basic facts about generalized $N$-functions and related Musielak–Orlicz Sobolev spaces. Further, we prove some auxiliary results and give the precise definition of the Nehari manifold to problem (${\rm P}_{\lambda}$) including its splitting into three disjoint parts. Section 3 discusses some basic estimates which are needed in the sequel while Section 4 gives a detailed study of the Nehari manifold and its properties. Finally, in Section 5, we are able to prove Theorem 1.1.
@@ 2. Preliminaries
Let $M(\Omega)$ be the set of all measurable functions $u\colon\Omega\to\mathbb{R}$. Then, the Musielak–Orlicz space $L^{\varphi}(\Omega)$ is defined by
@@
    Furthermore, if $u\in L^{\varphi}(\Omega)$, then $\rho_{\varphi}(u)<1$ (resp. =1=1; >1)>1) if and only if $\|u\|_{\varphi}<1$ (resp. =1=1; >1>1).
@@
We call a function $\varphi\colon\Omega\times[0,\infty)\to[0,\infty)$ a generalized $N$-function if $\varphi(\cdot,t)$ is measurable for all $t\in[0,\infty)$ and $\varphi(x,\cdot)$ is a $N$-function for a.a. $x\in\Omega$. We denote the class of all generalized $N$-functions by $N(\Omega)$.
@@ 1. $\bullet$
Next, we introduce the following functions, useful to compare generalized $N$-functions with suitable power functions. To this end, for given $\infty<\alpha\leq\beta<+\infty$, we define
@@
$\infty<i_{k}:=\inf_{s>0}\frac{sk^{\prime}(s)}{k(s)}\leq\sup_{s>0}\frac{sk^{\prime}(s)}{k(s)}=:s_{k}<+\infty.$ (2.5)
@@
It is worth noticing that (2.4) is automatically satisfied when $i_{k}>-1$, due to (2.11). This is the case of the NN-function $K:=\mathcal{H}(x,\cdot)$ (since (HHL)(H(HL)iii) forces $i_{k}>0$) and of the singular term $k:=f$ (see (HHf)).
@@
Proposition 2.6 (applied with $k=\partial_{s}\mathcal{H}(x,\cdot)$), (HHL)(H(HL)i), (HHL)(H(HL)iii), and (2.8) entail
@@
[…]
@@
$M(s)=\int_{0}^{s}m(t)\,{\rm d}t\quad\text{for all }s\in\mathbb{R}$
@@
$\mathcal{N}^{+}$ $:=\{u\in\mathcal{N}\colon\psi_{u}^{\prime\prime}(1)>0\},$
$\mathcal{N}^{0}$ $:=\{u\in\mathcal{N}\colon\psi_{u}^{\prime\prime}(1)=0\},$
$\mathcal{N}^{-}$ $:=\{u\in\mathcal{N}\colon\psi_{u}^{\prime\prime}(1)<0\}.$
@@
$E^{+}_{u}$ $:=\{t\in(0,+\infty)\colon tu\in\mathcal{N}^{+}\},$
$E^{0}_{u}$ $:=\{t\in(0,+\infty)\colon tu\in\mathcal{N}^{0}\},$
$E^{-}_{u}$ $:=\{t\in(0,+\infty)\colon tu\in\mathcal{N}^{-}\}.$
@@ 3. $\rm(a_{3})$}
where $\underline{W}_{p}^{q\theta},\overline{W}_{p}^{q\theta}$ are as in (2.3), while $\theta$ is defined in (2.13).
@@
Proposition 2.6, (HHL), and (2.13) yield
@@ 2. $\rm(b_{2})$}
In particular, $G$ is a generalized $N$-function satisfying the hypotheses of Proposition 2.8. Moreover, owing to (3.1), one has $\theta\geq 1$. Then (HHC) and Proposition 2.6 yield
@@ 4. Analysis of the Nehari manifold
with $c,C>0$, where we also used that $r_{-}>q\theta$, owing to (3.8). Recalling $1-\gamma_{-}<p$ yields $J(u_{n})\to+\infty$. ∎
@@
Thus, owing to Lemma 3.1, (HHf), (3.8), Proposition 2.6, and Lemma 3.2, we have
@@
Exploiting Lemma 3.1, (HHg), (3.8), Proposition 2.6, and Lemma 3.2, we deduce
@@
Fix any $u\in W^{1,\mathcal{H}}_{0}(\Omega)\setminus\{0\}$. We are going to apply Proposition 2.10 to the function $\psi^{\prime}_{u}$. To this end, using Lebesgue’s dominated convergence theorem and Fatou’s lemma, along with (HHm), (HHL), (HHg), (HHf), and the symmetry of $f$, ensures
@@ 5. Proof of the main result
$0$ $\leq\liminf_{n\to\infty}\langle\phi^{\prime}(\overline{t}\nabla u_{n})-\phi^{\prime}(\overline{t}\nabla u),\nabla u_{n}-\nabla u\rangle$
@@
$0$ $\leq\limsup_{n\to\infty}\frac{J(u+t_{n}h)-J(u)}{t_{n}}$ (5.5)
@@
$0$ $\leq\frac{J(\zeta(t_{n}h)(u+t_{n}h))-J(u)}{t_{n}}\leq\frac{J(\zeta(t_{n}h)(u+t_{n}h))-J(\zeta(t_{n}h)u)}{t_{n}}$ (5.6)
@@
$0$ $\leq\limsup_{n\to\infty}\frac{J(\zeta(t_{n}h)(u+t_{n}h))-J(u)}{t_{n}}$
@@
$0$ $\leq\langle T,(u+\varepsilon\varphi)^{+}\rangle$
@@
Notice that $|\{u+\varepsilon\varphi\leq 0\}|\to 0$ as $\varepsilon\to 0$. Therefore, $\langle T,\varphi\rangle\geq 0\quad\text{for all }\varphi\in W^{1,\mathcal{H}}_{0}(\Omega)$. Since $\varphi$ is arbitrarily chosen, we have $\langle T,\varphi\rangle=0$ for all $\varphi\in W^{1,\mathcal{H}}_{0}(\Omega)$, which means that $u$ is a weak solution to (${\rm P}_{\lambda}$). ∎